# Finite Volume Analysis of the Poisson Problem via a Reduced Discontinuous Galerkin Space

### v1

@@ Abstract
In this paper, we propose and analyze a high-order finite volume method for the Poisson problem based on the reduced discontinuous Galerkin (RDG) space. The main idea is to employ the RDG space as the trial space and the piecewise constant space as the test space, thereby formulating the scheme in a Petrov-Galerkin framework. This approach inherits the local conservation property of finite volume methods while benefiting from the approximation capabilities of discontinuous Galerkin spaces with significantly fewer degrees of freedom. We establish a rigorous error analysis of the proposed scheme: in particular, we prove optimal-order convergence in the DG energy norm and suboptimal-order convergence in $L^{2}$ norm. The theoretical analysis is supported by a set of one- and two-dimensional numerical experiments with Dirichlet and periodic boundary conditions, which confirm both the accuracy and efficiency of the method. The significance of this work lies in bridging finite volume and discontinuous Galerkin methodologies through the RDG space, thus enabling finite volume schemes with a mathematically rigorous convergence theory.
@@ 1 Introduction
The finite volume method (FVM) has a long and successful history in computational mathematics, renowned for its local conservation properties and its flexibility for handling complex geometries. Originally developed for hyperbolic conservation laws, notably by Godunov [godunov1959fvm] and later refined by van Leer [vanLeer1979fvm], FVM has also been extended to elliptic and parabolic problems such as the Poisson equation. Early developments include the SIMPLE algorithm of Patankar and Spalding [patankar1972fvm] for elliptic pressure equations, and subsequent gradient reconstruction techniques introduced in [patankar1980fvm]. While classical FVMs are effective, they often suffer from low-order accuracy, especially on distorted or skewed grids.
To improve accuracy, two strategies have been pursued. One is to employ high-order interface fluxes, such as ENO and WENO reconstructions [harten1987uniformly, liu1994weno]. Another is to approximate the numerical solution with higher-order polynomials. The discontinuous Galerkin (DG) method [cockburn2001discontinuous] is a prominent representative of the latter approach, offering high-order accuracy and geometric flexibility. However, standard DG methods involve large numbers of degrees of freedom, which leads to high computational cost. To address this, reconstruction-based DG methods have been proposed. For example, Li et al. [liruo2019arbitrary, liruo2019biharmonic, liruo2020elasticity] introduced patch-reconstruction techniques that dramatically reduce the number of unknowns while retaining high-order accuracy. Similar ideas can be traced back to the finite element literature, e.g. [onate1993derivation, hampshire1992three].
Recently, Hou and Xia [hou2024rdg] constructed a reduced discontinuous Galerkin (RDG) space, which can achieve the same approximation order as the standard DG space but with fewer degrees of freedom. This motivates the present study: we aim to leverage the RDG space to design and analyze finite volume schemes for elliptic problems.
Compared to DG and FEM, the FVM framework traditionally lacks a unified variational foundation, which makes rigorous error analysis more fragmented. Several efforts have been made to close this gap. Süli [suli1991poisson] interpreted cell-centered finite volume schemes for the Poisson equation as Petrov-Galerkin methods, obtaining optimal $H^{1}$-error estimates and suboptimal $L^{\infty}$-error estimates. Extensions and refinements include the works of Jones et al. [jones2000fvm], Sheng et al. [yuan2007fvm], Hermeline [hermeline2000fvm], and Eymard et al. [droniou2014fvm]. Nevertheless, compared with DG analysis, the finite volume theory remains less unified, particularly in higher-order settings.
The main goal of this paper is to apply the reduced DG space introduced in [hou2024rdg] to formulate and analyze a finite volume scheme for the Poisson model problem. Specifically, we adopt the RDG space as the trial space and the piecewise constant space as the test space, leading to a Petrov-Galerkin formulation within the symmetric interior penalty DG framework. We establish approximation properties, derive error estimates, and validate the method through numerical experiments. The proposed method achieves optimal-order convergence in the DG energy norm and suboptimal-order convergence in the $L^{2}$ norm, as confirmed by numerical tests in one and two dimensions with Dirichlet and periodic boundary conditions.
@@
The rest of this paper is organized as follows. In Section 2, we introduce the notations, definitions, and preliminaries. Section 3 recalls the RDG space and its approximation properties. Section 4 presents the proposed finite volume scheme for the Poisson problem, together with rigorous error estimates in the DG energy norm and $L^{2}$-norm. Section 5 reports numerical experiments that validate the theoretical findings. Finally, Section 6 concludes the paper with remarks and perspectives.
@@ 3.2 Properties of the reconstruction operator
This assumption guarantees the uniqueness of the reconstruction and will be assumed henceforth.
@@ 4.1 Dirichlet Problem
For the bilinear form $a_{h}^{\mathrm{sip}}:U^{2}_{h}\times V^{0}_{h}\to\mathbb{R}$, when $\eta\in(-\frac{3}{2},5)$, the inf–sup condition holds:
@@
The above analysis ensures that $C_{j}>0$ for $\tfrac{3}{2}<\eta<5$, thereby guaranteeing the coercivity of $a_{h}^{\mathrm{SIP}}(\cdot,\cdot)$. Nevertheless, numerical experiments suggest that this coercivity property may remain valid for a wider range of $\eta$. It is worth noting that this requirement differs from that of the classical IPDG scheme, where the penalty parameter $\eta>0$ must be sufficiently large, depending on the discrete trace inequality.
@@
Moreover, by the definition of the discrete problem,
@@
$=\sum_{i,j}\left(\mathcal{R}^{(1)}_{i,j}+\mathcal{R}^{(2)}_{i,j}+\mathcal{R}^{(3)}_{i,j}+\mathcal{R}^{(4)}_{i,j}\right)=\mathcal{R}_{1}+\mathcal{R}_{2}+\mathcal{R}_{3}+\mathcal{R}_{4}.$
@@
$\mathcal{R}^{(1)}_{i,j}$ $=\int_{x_{i}}^{x_{i+1}}\int_{y_{j}}^{y_{j+1}}u_{x}^{2}\,\mathrm{d}x\,\mathrm{d}y,$
$\mathcal{R}^{(2)}_{i,j}$ $=\int_{x_{i}}^{x_{i+1}}\int_{y_{j}}^{y_{j+1}}u_{y}^{2}\,\mathrm{d}x\,\mathrm{d}y,$
@@
$:=\sum_{i,j}\big(\mathcal{R}^{(5)}_{i,j}+\mathcal{R}^{(6)}_{i,j}\big):=\mathcal{R}_{5}+\mathcal{R}_{6},$
@@
$\quad\begin{aligned} A_{i,j}&:=-\tfrac{1}{2}D_{(i,j-1/2)}u+\tfrac{1}{2}D_{(i+1,j-1/2)}u+D_{(i,j+1/2)}u-D_{(i+1,j+1/2)}u\\[2.0pt] &\qquad\quad-\tfrac{1}{2}D_{(i,j+3/2)}u+\tfrac{1}{2}D_{(i+1,j+3/2)}u,\\ B_{i,j}&:=D_{(i,j-1/2)}u-2D_{(i,j+1/2)}u+D_{(i,j+3/2)}u,\\ C_{i,j}&:=\tfrac{1}{2}D_{(i,j-1/2)}u-\tfrac{1}{2}D_{(i+1,j-1/2)}u-\tfrac{1}{2}D_{(i,j+1/2)}u+\tfrac{1}{2}D_{(i+1,j+1/2)}u,\\ D_{i,j}&:=-\tfrac{1}{6}D_{(i-1/2,j)}u+\tfrac{1}{6}D_{(i-1/2,j+1)}u-\tfrac{5}{12}D_{(i+1/2,j)}u+\tfrac{5}{12}D_{(i+1/2,j+1)}u\\[2.0pt] &\qquad\quad+\tfrac{1}{12}D_{(i+3/2,j)}u-\tfrac{1}{12}D_{(i+3/2,j+1)}u,\\ E_{i,j}&:=-D_{(i,j-1/2)}u+D_{(i,j+1/2)}u,\\ F_{i,j}&:=\tfrac{1}{3}D_{(i-1/2,j)}u+\tfrac{5}{6}D_{(i+1/2,j)}u-\tfrac{1}{6}D_{(i+3/2,j)}u.\end{aligned}$
@@
where $\bm{M}\in\mathbb{R}^{8\times 8}$ is symmetric positive definite whose eigenvalues are bounded above and below by positive constants independent of the mesh size. (see Appendix C for details).
@@
For the bilinear form $a^{\mathrm{sip}}_{h}:U^{2}_{h}\times V^{0}_{h}\to\mathbb{R}$, when $\eta\in(-\frac{3}{2},5)$, the following inf–sup condition holds:
@@ 4.2 Periodic Problem
To formulate the problem, we introduce a Lagrange multiplier to enforce the mean-zero condition. The corresponding bilinear form $B(\cdot,\cdot)$ is defined for all $u,v_{h}\in H^{2}(\mathcal{T}_{h})$ and $\lambda,\mu\in\mathbb{R}$ by
$B((u,\lambda),(v_{h},\mu))$ $=a^{0}_{h}(u,v_{h})+b(\lambda,v_{h})+b(u,\mu),$
@@
The periodic boundary condition can be naturally enforced within the virtual elements $\mathcal{T}_{h}^{\text{virt}}$ and extended to higher-order cases with larger virtual stencils. Therefore, the analysis for the periodic problem can be readily generalized to even values of $k$. The equivalence between the IPDG scheme (4.24) and the finite volume method follows from Lemmas 4.1 and 4.8, which implies that the second and third terms in the equivalent form of $a^{0}_{h}(\cdot,\cdot)$ vanish.
@@ 5.1 Convergence Study for the Dirichlet Problem
Another interesting observation is that the finite volume scheme remains stable and accurate even for $\eta=0$ or negative values such as $\eta=-1.4$. This observation is consistent with the theoretical results in Lemmas 4.5 and 4.11, which establish that the inf-sup condition holds for $\eta\in(-\tfrac{3}{2},5)$. Moreover, when $\eta=0$, the scheme exhibits better accuracy compared to cases with $\eta\neq 0$. In this special case, the scheme (4.4) reduces exactly to the classical second-order finite volume scheme for the Poisson problem. This further highlights the distinction between the penalty parameter requirements of the finite volume (Petrov-Galerkin) formulation and those of the classical IPDG method.
@@ 6 Conclusions
$\frac{g_{k}^{\prime}(x)}{g_{k}(x)}=\frac{L^{\prime}_{k}(x)}{L_{k}(x)}-\frac{1}{x}=\sum_{\begin{subarray}{c}r\in\{-m,\dots,m+1\}\\ r\neq 0,\,r\neq k\end{subarray}}\frac{1}{x-r}:=H_{k}(x).$ (A.5)
@@
$g_{k}(0)=\frac{\prod_{r=-m,\,r\neq 0,\,k}^{m+1}(-r)}{\prod_{r=-m,\,r\neq k}^{m+1}(k-r)}.$
@@
The inductive step proceeds in the same manner as in the odd-$m$ case, thereby completing the proof for even $m$.
@@
We prove the result for the boundary face $\{x_{i}\}\times[y_{j},y_{j+1}]$; the proofs for other faces are analogous and thus omitted for brevity. Let $\{\phi_{p}(x)\}_{p=-m}^{m}$ and $\{\phi_{q}(y)\}_{q=-m}^{m}$ denote the set of one-dimensional basis functions on the reduced Galerkin space. For each element $K_{i,j}=[x_{j},x_{j+1}]\times[y_{i},y_{i+1}]$, there exists an affine mapping from the reference element $\hat{K}=[0,1]^{2}$. This mapping, $F_{i,j}:\hat{K}\to K_{i,j}$, transforms the reference coordinates $\hat{x}=(\hat{x},\hat{y})\in\hat{K}$ to the physical coordinate $x=(x,y)\in K_{i,j}$ as follow:

### v2

@@ Abstract
In this paper, we analyze the finite volume method for the Poisson equation by recasting it as a discontinuous Petrov-Galerkin scheme built upon the reduced discontinuous Galerkin (RDG) space. The key idea is to employ the RDG space as the trial space and piecewise constants as the test space, thereby formulating the method within a Petrov-Galerkin framework. This construction preserves the local conservation property characteristic of finite volume methods, while leveraging the approximation power of discontinuous Galerkin spaces at substantially reduced computational cost. An interesting observation is that the penalty parameter arising in this formulation differs markedly from its counterpart in standard discontinuous Galerkin methods. We provide a rigorous error analysis of the proposed scheme. In particular, we establish optimal-order convergence in the DG energy norm and suboptimal-order convergence rate in the $L^{2}$ norm. The theoretical analysis is supported by a set of one- and two-dimensional numerical experiments with Dirichlet and periodic boundary conditions. This study highlights the role of the RDG space as an effective bridge between finite volume and discontinuous Galerkin methodologies, enabling finite volume schemes equipped with a mathematically rigorous convergence theory.
@@ 1 Introduction
The finite volume method (FVM) has long been a cornerstone in computational fluid dynamics and heat transfer simulations due to its inherent local mass conservation and geometric flexibility [godunov1959fvm, patankar1972fvm, patankar1980fvm, eymard2000fvm, vanLeer1979fvm, harten1987uniformly, liu1994weno, onate1993derivation, hampshire1992three]. Despite its widespread engineering success, establishing a rigorous theoretical framework for FVM has historically presented significant challenges. Unlike the finite element method (FEM), which rests on a clear variational foundation, the classical FVM was originally derived directly from integral conservation laws. Consequently, early error analyses were often dependent on geometric constraints, such as requiring admissible meshes where the line connecting cell centers is orthogonal to the interface. To overcome these limitations and establish a sound mathematical framework, theoretical developments have generally evolved along two distinct paths: vertex-centered and cell-centered formulations.
Research on vertex-centered schemes which are often referred to as finite volume element methods, is relatively mature, largely benefiting from their close relationship with linear finite element methods. Seminal works in the 1980s and 1990s laid the foundation. Bank and Rose [bank1987fvm] established the rigorous theoretical framework for the vertex-centered box method applied to elliptic problems by interpreting it as a Petrov-Galerkin scheme. They demonstrated that, for the Poisson equation, the stiffness matrix is identical to that of linear FEM, thereby guaranteeing optimal convergence in the energy norm, comparable to that of standard linear finite elements. Subsequently, Cai [cai1991fvem] systematized this theory for general self-adjoint elliptic equations, deriving $O(h)$ error estimates in the discrete $H^{1}$ norm on shape-regular unstructured triangulations. Regarding the geometric constraints on triangular meshes, Mishev [mishev1998fvm] demonstrated that the classic vertex-centered scheme for elliptic problems yields optimal $O(h)$ convergence and preserves monotonicity, provided that the underlying mesh satisfies the Voronoi property. More recently, efforts have been extended to high-order formulations, where significant progress has been made through different theoretical approaches. Ye [ye2004fvm] proposed a discontinuous finite volume (DFV) method for elliptic problems, using discontinuous piecewise polynomials and penalty terms to ensure stability on minimal-support dual partitions, thereby establishing optimal $O(h)$ convergence in the discrete $H^{1}$ norm. Building on this, Chou and Ye [ye2007fvm] developed a unified analysis framework covering conforming, nonconforming, and discontinuous finite volume schemes. Through a rigorous duality argument within this framework, they successfully derived the previously elusive optimal $O(h^{2})$ $L^{2}$ norm error estimates. Building on the algebraic connections established by Bank and Rose, Xu and Zou [xu2009fvm] derived a generalized identity between the stiffness matrices of FVM and FEM. They established the inf-sup condition for both linear and quadratic vertex-centered schemes on simplicial meshes, thereby extending the optimal convergence theory to higher-order approximations. Chen [chen2010fvm] simplified the geometric construction by introducing a hierarchical decomposition strategy, applying finite volume constraints to vertex unknowns and Galerkin projections to high-order bubble functions. This hybrid structure ensures the inf-sup condition on unstructured meshes without requiring complex dual partitions for non-vertex degrees of freedom. Cao et al. [cao2013fvm] focused on enhancing accuracy by exploiting superconvergence. By defining trial functions on Gauss-Lobatto nodes and constructing dual partition boundaries at Gauss points, they achieved superior convergence rates at nodal points compared to standard global estimates.
Conversely, the theoretical landscape for cell-centered formulations, which constitutes the focus of this work, has evolved through a fundamentally different path, primarily relying on connections to the finite difference method (FD), the equivalence with mixed finite element methods and Petrov-Galerkin variational formulations. Forsyth et al. [forsyth1988fvm] utilized maximum principles to establish quadratic convergence for cell-centered schemes on smooth grids, effectively resolving earlier concerns regarding local truncation errors. Jones and Menzies [jones2000fvm] addressed the inconsistency on nonuniform rectangular grids. They demonstrated that, although the local truncation error degrades to $O(1)$ due to mesh irregularity, the stability of the equivalent finite difference operator ensures $O(h^{2})$ global convergence. Süli [suli1991poisson] formulated the scheme as a Petrov-Galerkin method, establishing optimal $O(h)$ convergence in the discrete $H^{1}$ norm on nonuniform meshes, while also demonstrating $O(h^{2})$ convergence under higher regularity assumptions and mild mesh constraints. Parallel to this, Chavent and Roberts [chavent1991fvm] and Arbogast et al. [arbogast1997fvm] demonstrated that, specifically on rectangular grids, cell-centered finite volume schemes are algebraically equivalent to the lowest-order Raviart-Thomas ($RT_{0}$) mixed method with specific quadrature rules. Subsequently, Arbogast et al. [arbogast1998fvm] significantly extended this framework to geometrically general domains. They derived cell-centered schemes for full tensor coefficients on irregular meshes by utilizing an expanded mixed formulation. Although they introduced an enhanced technique with interface Lagrange multipliers to handle mesh distortions, their analysis remained confined to the lowest-order approximation. Extending this variational equivalence to high-order schemes on general unstructured triangulations remains open. On rectangular grids, however, substantial progress has been made. Cai et al. [cai2003fvm] systematically derived high-order finite volume schemes based on $RTN$ and $BDFM$ mixed spaces. Crucially, they acknowledged the intrinsic limitations of this approach. For high-order $BDFM$ spaces, a satisfactory finite volume reduction is currently unachievable. Furthermore, they noted that as the polynomial degree increases, the computational efficiency advantage of the derived FVM diminishes, observing that the resulting finite volume scheme may actually require more degrees of freedom than the underlying hybridized mixed finite element method.
Subsequently, theoretical research diverged. While unified frameworks like the gradient discretisation method [droniou2018gdm] solidified the analysis for low-order schemes on general meshes, the pursuit of high-order accuracy increasingly converged towards DG methods. However, despite their rigorous variational foundation, DG methods typically incur high computational costs due to the large number of coupled degrees of freedom.
In this paper, we address this challenge by formulating the cell-centered finite volume scheme as a variational Petrov-Galerkin method to analyze its convergence properties for the Poisson model problem on rectangular grids. Specifically, we utilize the reduced discontinuous Galerkin (RDG) space introduced in [hou2024rdg] as the trial space, while retaining the piecewise constant space as the test space. This formulation constitutes a high-order FVM. We rigorously prove that, on uniform meshes, this variational formulation is mathematically equivalent to the classical cell-centered FVM. This equivalence allows us to leverage the theories of DG methods to derive optimal error estimates in the DG energy norm and suboptimal error estimates in the $L^{2}$ norm. Validated by numerical tests, this work provides a rigorous theoretical foundation for high-order FVMs on structured grids.
@@
The rest of this paper is organized as follows. In Section 2, we introduce the notations, definitions, and preliminaries. Section 3 recalls the RDG space and its approximation properties. Section 4 presents the proposed finite volume scheme for the Poisson problem, together with rigorous error estimates in the DG energy norm and $L^{2}$ norm. Section 5 reports numerical experiments that validate the theoretical findings. Finally, Section 6 concludes the paper with remarks and perspectives.
@@ 3.2 Properties of the reconstruction operator
This assumption guarantees uniqueness of the reconstruction and will be assumed henceforth.
@@ 4.1 Dirichlet Problem
For the bilinear form $a_{h}^{\mathrm{sip}}:U^{2}_{h}\times V^{0}_{h}\to\mathbb{R}$, the inf–sup condition holds:
@@
The above analysis ensures $C_{j}>0$ for $\tfrac{3}{2}<\eta<5$, which guarantees the coercivity of $a_{h}^{\mathrm{sip}}(\cdot,\cdot)$. Nevertheless, numerical experiments indicate that this property may hold for a wider range of $\eta$.
@@
Moreover, by definition of the discrete problem,
@@
$=\sum_{i,j}\left(\mathcal{R}^{(1)}_{i,j}+\mathcal{R}^{(2)}_{i,j}+\mathcal{R}^{(3)}_{i,j}+\mathcal{R}^{(4)}_{i,j}\right)\coloneq\mathcal{R}_{1}+\mathcal{R}_{2}+\mathcal{R}_{3}+\mathcal{R}_{4}.$
@@
$\mathcal{R}^{(1)}_{i,j}$ $\coloneq\int_{x_{i}}^{x_{i+1}}\int_{y_{j}}^{y_{j+1}}u_{x}^{2}\,\mathrm{d}x\,\mathrm{d}y,$
$\mathcal{R}^{(2)}_{i,j}$ $\coloneq\int_{x_{i}}^{x_{i+1}}\int_{y_{j}}^{y_{j+1}}u_{y}^{2}\,\mathrm{d}x\,\mathrm{d}y,$
@@
$\coloneq\sum_{i,j}\big(\mathcal{R}^{(5)}_{i,j}+\mathcal{R}^{(6)}_{i,j}\big)\coloneq\mathcal{R}_{5}+\mathcal{R}_{6},$
@@
$\quad\begin{aligned} A_{i,j}&\coloneq-\tfrac{1}{2}D_{(i,j-1/2)}u+\tfrac{1}{2}D_{(i+1,j-1/2)}u+D_{(i,j+1/2)}u-D_{(i+1,j+1/2)}u\\[2.0pt] &\qquad\quad-\tfrac{1}{2}D_{(i,j+3/2)}u+\tfrac{1}{2}D_{(i+1,j+3/2)}u,\\ B_{i,j}&\coloneq D_{(i,j-1/2)}u-2D_{(i,j+1/2)}u+D_{(i,j+3/2)}u,\\ C_{i,j}&\coloneq\tfrac{1}{2}D_{(i,j-1/2)}u-\tfrac{1}{2}D_{(i+1,j-1/2)}u-\tfrac{1}{2}D_{(i,j+1/2)}u+\tfrac{1}{2}D_{(i+1,j+1/2)}u,\\ D_{i,j}&\coloneq-\tfrac{1}{6}D_{(i-1/2,j)}u+\tfrac{1}{6}D_{(i-1/2,j+1)}u-\tfrac{5}{12}D_{(i+1/2,j)}u+\tfrac{5}{12}D_{(i+1/2,j+1)}u\\[2.0pt] &\qquad\quad+\tfrac{1}{12}D_{(i+3/2,j)}u-\tfrac{1}{12}D_{(i+3/2,j+1)}u,\\ E_{i,j}&\coloneq-D_{(i,j-1/2)}u+D_{(i,j+1/2)}u,\\ F_{i,j}&\coloneq\tfrac{1}{3}D_{(i-1/2,j)}u+\tfrac{5}{6}D_{(i+1/2,j)}u-\tfrac{1}{6}D_{(i+3/2,j)}u.\end{aligned}$
@@
where $\bm{M}\in\mathbb{R}^{8\times 8}$ is symmetric positive define whose eigenvalues are bounded above and below by positive constants independent of the mesh size. (see Appendix C for details).
@@
For the bilinear form $a^{\mathrm{sip}}_{h}:U^{2}_{h}\times V^{0}_{h}\to\mathbb{R}$, the following inf–sup condition holds:
@@ 4.2 Periodic Problem
To formulate the problem, we introduce a Lagrange multiplier to enforce the mean-zero condition. The corresponding bilinear form $B(\cdot,\cdot)$ is defined for all $u,v\in H^{1}(\Omega,\mathcal{T}_{h})$ and $\lambda,\mu\in\mathbb{R}$ by
$B((u,\lambda),(v,\mu))$ $=a^{0}_{h}(u,v)+b(\lambda,v)+b(u,\mu),$
@@
The periodic boundary condition can be naturally enforced within the virtual elements $\mathcal{T}_{h}^{\text{virt}}$ and extended to higher-order cases with larger virtual stencils. Therefore, the analysis for the periodic problem can be readily generalized to even values of $k$. The equivalence between the IPDG scheme (4.24) and the finite volume method follows from Lemmas 4.1 and 4.8, which implies that the second and third terms in (4.24) vanish.
@@ 6 Conclusions
$\frac{g_{k}^{\prime}(x)}{g_{k}(x)}=\frac{L^{\prime}_{k}(x)}{L_{k}(x)}-\frac{1}{x}=\sum_{\begin{subarray}{c}r\in\{-m,\dots,m+1\}\\ r\neq 0,\,r\neq k\end{subarray}}\frac{1}{x-r}\coloneq H_{k}(x).$ (A.5)
@@
$g_{k}(0)=\frac{\prod_{r=-m,\,r\neq 0,\,k}^{m+1}(-r)}{\prod_{r=-m,\,r\neq k}^{m+1}(k-r)},$
@@
The inductive step proceeds in exactly the same manner as in the odd-$m$ case, thereby completing the proof for even $m$.
@@
We prove the result for the boundary face $\{x_{i}\}\times[y_{j},y_{j+1}]$; the proofs for other faces are analogous and thus omitted for brevity. Let $\{\phi_{p}(x)\}_{p=-m}^{m}$ and $\{\phi_{q}(y)\}_{q=-m}^{m}$ denote the set of one-dimensional basis functions on reduced Galerkin space. For each element $K_{i,j}=[x_{j},x_{j+1}]\times[y_{i},y_{i+1}]$, there exists an affine mapping from the reference element $\hat{K}=[0,1]^{2}$. This mapping, $F_{i,j}:\hat{K}\to K_{i,j}$, transforms the reference coordinates $\hat{x}=(\hat{x},\hat{y})\in\hat{K}$ to the physical coordinate $x=(x,y)\in K_{i,j}$ as follow: